84=xsquared -17x+66 what are the possible values of x? with steps please
Answer by stanbon(75887)

84=xsquared -17x+66 what are the possible values of x? with steps please
----
x^2 - 17x + 66 = 84
---
x^2 - 17x - 18 = 0
Factor:
(x-18)(x+1) = 0
x = 18 or x = -1
=========================
Cheers,
Stan H.